# Learning Object State Changes in Videos: An Open-World Perspective

### v1

@@ Abstract
Object State Changes (OSCs) are pivotal for video understanding. While humans can effortlessly generalize OSC understanding from familiar to unknown objects, current approaches are confined to a closed vocabulary. Addressing this gap, we introduce a novel open-world formulation for the video OSC problem. The goal is to temporally localize the three stages of an OSC—the object’s initial state, its transitioning state, and its end state—whether or not the object has been observed during training. Towards this end, we develop V ID OSC, a holistic learning approach that: (1) leverages text and vision-language models for supervisory signals to obviate manually labeling OSC training data, and (2) abstracts fine-grained shared state representations from objects to enhance generalization. Furthermore, we present HowToChange, the first open-world benchmark for video OSC localization, which offers an order of magnitude increase in the label space and annotation volume compared to the best existing benchmark. Experimental results demonstrate the efficacy of our approach, in both traditional closed-world and open-world scenarios.
@@ 1 Introduction
However, the development of video OSC is still at a primitive stage. Existing approaches [2, 49, 50, 14, 1, 45, 57] assume a closed vocabulary, where each state transformation is associated with a limited set of objects—often just 1 or 2. For example, the concept of “melting” might be limited to familiar items like butter and marshmallow. Consequently, the learned models are only capable of identifying state changes for objects observed during training and stumble when presented with novel objects. In contrast, in real-world situations a single state transition like “melting” can be linked with a plethora of objects—some ubiquitous and others rare. Importantly, there is an intrinsic connection that threads these OSC processes together. As humans, even if we have never encountered an unusual substance (e.g., jaggery) before, we can still understand that it is experiencing a “melting” process based on its visual transformation cues (see Fig. 1, bottom).
@@
Armed with this new problem formulation, we propose a holistic video OSC learning approach, anchored by two innovative ideas. First, we explore text and vision-language models (VLMs) for supervisory signals during training. We pioneer the use of textual state descriptions to generate a long tail of object state pseudo-labels, by leveraging the remarkable capabilities of VLMs [43, 22, 62] and large language models (LLMs) [52, 53, 39]. This strategy eliminates the need for exhaustive label collection for training data, facilitating large-scale model training. [^1] Note that at test time, our model requires only the video, with no need for additional text, ensuring utmost flexibility and applicability. Second, to confront the open-world challenge, we propose object-agnostic state prediction, consisting of three key techniques: a shared state vocabulary to unify the label space across objects sharing the same state transition (e.g., melting butter and melting marshmallow), temporal modeling to grasp the progression of state changes over time, and object-centric features that better represent the objects during an OSC. These designs equip our model with the ability to generalize state understanding observed in known objects to novel ones. We term our approach VidOSC.
@@ 2 Related Work
Object State Changes Image-based methods explore the compositional nature of objects and their attributes, including zero-shot recognition of unseen combinations [36, 38, 21, 42, 32, 37, 40, 28], but do not consider the temporal progression of OSCs in video, which brings new challenges. Video-based methods develop human-centric models that leverage state change cues to facilitate action recognition  [14, 57, 1, 45], or explore joint discovery of object states and actions  [2, 49, 50]. Notably, all the existing methods assume a closed world, recognizing only the OSC categories (“known” objects) seen during training. Our approach distinguishes itself in three crucial ways: (1) we introduce a novel open-world formulation, where the objects encountered during evaluation can be entirely unseen during training; (2) we adopt an object-centric perspective, allowing for scenarios where OSCs occur with no observable human actions in the video; and (3) we propose to utilize the text modality and VLMs as supervisory signals, greatly scaling up model training and boosting performance.
@@
Vision and Language LLMs [52, 53, 39] have revolutionized various research fields with their exceptional performance across diverse applications. Building on this momentum, the use of web-scale image-text data has emerged as a powerful paradigm in computer vision, and led to powerful VLMs that advance an array of image tasks, including zero-shot classification [43], detection [17], segmentation [31] and visual question answering [26, 27]. Similarly, joint video-text representation learning [35, 29, 66, 3] has been advanced by multimodal datasets like HowTo100M [35] and Ego4D [16], which offer large-scale collections of videos paired with their corresponding narrations. These datasets have also facilitated advances like step discovery [11], localization in procedural activities [34] and long egocentric videos [44]. In line with this, we propose to leverage the text accompanying instructional videos as well as existing high-performing VLMs to provide supervisory signals that guide the training of our video OSC model and allow learning for the open world.
@@ 3 Approach
Figure 2: Our proposed VidOSC framework: (a) Mining for OSC examples (Sec. 3.3): We leverage ASR transcriptions paired with videos and the capabilities of LLM to automatically mine OSC examples. (b) Pseudo Label Generation (Sec. 3.3): We utilize textual state descriptions and an VLM for supervisory signals during training; (c) Model Training (Sec. 3.2): We develop a video model for object-agnostic state prediction. (d) Model Testing (Sec. 3.1): We propose an open-world formulation, evaluating on both known and novel OSCs. Notably, while we employ the text modality to guide model training, our model is purely video-based and requires no text input at the test phase, ensuring maximum flexibility and applicability. Ground truth for the test set is manually annotated.
@@ 3.1 Video OSC in the Open World
We begin by formally defining an object state change (OSC) in videos as a visually detectable transformation, where an object experiences a change that is not easily reversible, in line with [16]. We characterize an OSC category as an object combined with a specific state transition, such as “chicken + shredding”, and delineate an OSC process through three distinct states: initial (precondition state), transitioning, and end (postcondition state). [^2] Due to real-world data variation, some videos may lack one of the OSC states, and there may be multiple segments in a video corresponding to the same state. Our formulation accounts for all these scenarios. It is essential to highlight that we take an object-centric perspective: the “transitioning state” accounts for instances where the video depicts an active action applied to the object, as well as scenarios where the object undergoes a passive transformation absent of human intervention.
@@
Next, we propose an open-world problem formulation, capturing the intrinsic long-tail distribution observed in real-world scenarios. Consider $N$ state transitions, each paired with a set of associated objects, denoted by $\mathbf{O}^{n}=\{o_{1}^{n},o_{2}^{n},\cdots,o_{m(n)}^{n}\}$, for $n=\{1,2,\cdots,N\}$, where $m(n)$ denotes the number of objects linked with the $n$-th state transition. Within a specific state transition (like frying), certain objects in the set (such as chicken or onions) are frequently observed, while the combination of the same state transition with other objects (such as cauliflower) appear less often. Motivated by this inherent long-tail, we propose to split $\mathbf{O}^{n}$ into two distinct subsets: $\mathbf{O}_{\text{known}}^{n}$ covering the common combinations observed during training, and $\mathbf{O}_{\text{novel}}^{n}$ comprising the infrequent combinations, which are unseen during training due to their rarity. During test, the model’s predictions are evaluated on the entire object set $\mathbf{O}^{n}$ for a comprehensive reflection of the open-world setting.
@@ 3.2 Object-Agnostic State Prediction
Shared State Vocabulary The inherent link among different OSC categories is overlooked in prior work. Common practice involves developing separate models for each OSC category [2, 49] (e.g., one model exclusively for melting butter and another for melting chocolate), or using one single model that still treats every OSC as a distinct entity in the label space [50] (e.g., considering melted butter and melted chocolate as two separate end states). Such a formulation results in an extensive label set with $K=3\times\sum_{n=1}^{N}m(n)$, where 3 denotes the number of OSC states (i.e., initial, transitioning and end) and $\sum_{n=1}^{N}m(n)$ is the total number of OSC categories. This compartmentalized view can inadvertently hinder the model’s generalization ability. Yet, on closer inspection, states across varied objects are intrinsically related. For instance, the “melting” principle remains consistent, even when applied to visually distinct objects like butter and chocolate. This motivates us to group training of OSCs that share the same state transition together, and treat same-state labels across objects as one to reduce the label set size to $K<3\times\sum_{n=1}^{N}m(n)$. Consequently, the model is encouraged to learn common representations that characterize the initial, transitioning, and end states of “melting” that are shared by both butter and chocolate. This state sharing strategy not only simplifies training but also aids the model in transferring knowledge of state transitions from known to novel objects.
@@
Temporal Modeling Recognizing that state changes unfold over time, it is important to capture the temporal dynamics in videos. An object’s state at any frame, be it initial, transitioning, or end, is often best understood in the context of preceding and succeeding frames. Contrary to prior works [2, 49, 50] that rely on isolated frame-wise modeling without considering the temporal progression of video OSCs, we address this gap by proposing a temporally-aware model design. As illustrated in Fig. 2(c), we adopt an ecnoder-decoder architecture. For the encoding phase, input video features $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{T}\}$ are first projected using an MLP $f_{\text{project}}$, and augmented with sinusoidal positional embeddings, $\mathbf{Z}_{\text{pos}}$. Subsequently, a transformer encoder [55] $f_{\text{transformer}}$ is adopted to capture the temporal dynamics among these features, yielding $\mathbf{Z}=f_{\text{transformer}}(f_{\text{project}}(\mathbf{X})+\mathbf{Z}_{\text{pos}})$. For the decoding phase, a MLP decoder $g$, maps these temporally-aware hidden representations to OSC state predictions: $\tilde{\mathbf{Y}}=g(\mathbf{Z})$. See Sec. 5 for architecture and training details. This design ensures that when predicting the state for a frame, the model has assimilated temporal context from the entire sequence. Essentially, our model exploits the fact that the dynamics of the state change has greater invariance to the object category than how the object looks, emphasizing object transformations along time over object static appearances, to enhance generalization capabilities for novel OSCs.
@@ 3.3 Text and VLMs as Supervision
Mining for OSC examples Utilizing a vast collection of how-to instrutional videos as the training source, we develop an automated mining process to capture the rich, real-world variability of OSCs. The motivation is that instructional videos usually have accompanying ASR transcriptions that offer valuable OSC cues. For example, a speaker mentioning, “so just get your rotisserie chicken and shred it up” suggests that the chicken may undergo a state transition of shredding in the video. Leveraging this finding, we employ LLAMA2 [53] to analyze ASR transcriptions for identifying candidate videos and their associated OSC categories. This text mining stage allows us to corral a long tail of OSCs, discovering likely state change terms—even if rare—from the free-form natural language narrations given by how-to instructors in the training videos. See Fig. 2 (a).
@@
$\hat{y}_{t}=\begin{cases}\text{Initial}&\text{if }\mathbf{S}[t,0]-\mathbf{S}[t,1]>\delta\text{ and }\mathbf{S}[t,0]-\mathbf{S}[t,2]>\delta\\ \text{Transitioning}&\text{if }\mathbf{S}[t,1]-\mathbf{S}[t,0]>\delta\text{ and }\mathbf{S}[t,1]-\mathbf{S}[t,2]>\delta\\ \text{End}&\text{if }\mathbf{S}[t,2]-\mathbf{S}[t,0]>\delta\text{ and }\mathbf{S}[t,2]-\mathbf{S}[t,1]>\delta\\ \text{Background}&\text{if }{\sum}(\mathbf{S}[t,:])<\tau\\ \text{Ambiguous}&\text{otherwise}\end{cases}$
@@
To further refine these labels, we enforce a causal ordering constraint. Given the inherent progression of OSCs, the anticipated order is initial states followed by transitioning states, and finally, the end states. Any frame whose pseudo label does not respect this natural progression is re-assigned to the ambiguous category. Our training employs a cross-entropy loss between pseudo label $\hat{y}_{t}$ and the corresponding model prediction $\tilde{y}_{t}$, with ambiguous frames excluded to maintain clarity and distinction among the state labels.
@@ 4 The HowToChange Dataset
Existing OSC datasets fall short in capturing the open-world’s diversity and long-tail distribution of OSCs. To address this gap, we introduce HowToChange. It encompasses varied state transitions coupled with a diverse range of objects, providing an authentic reflection of the real-world frequency, from the commonplace to the rare. Below, we detail our data and label collection process.
@@
Ground Truth Label Collection (Evaluation Set) To facilitate thorough evaluation, we obtain manual annotations for a subset of 5,424 videos from the larger dataset. [^4] Our training is purely guided by VLM and requires no ground truth labels. The annotations are reserved exclusively for evaluation. The annotation workflow is as follows: each annotator is presented with a video segment along with an OSC category that was previously identified by LLAMA2. The annotator has the option to reject the video segment if it does not contain the specified OSC. Otherwise, they label the time ranges corresponding to the initial, transitioning, and end states of the OSC. Adhering to our object-centric emphasis, annotators are instructed to label based on the visual changes of the object, rather than human-centric actions, and exclude time ranges where the object of interest is not visible, ensuring clean and focused temporal labels. Fig. 3 provides the distribution of annotated videos. On average, we collect 271 annotations per state transition, with a video duration of 41.5 seconds, and 12.9% of videos belong to the novel category. The entire annotation process required around 1,507 hours by 30 professional annotators.
 Datasets # Obj # ST # OSC Avg. # Videos GT Label?
@@
Table 1: Comparison with existing video datasets focusing on object states. ‘Obj’ and ‘ST’ represent objects and state transitions, respectively. ‘Avg.’ denotes the average number of objects associated with a state transition; higher values indicate more need to generalization across objects. We present the first open-world benchmark for temporal video OSC, with an order of magnitude increase in OSC vocabulary and annotation volume.
Dataset Comparison Table 1 compares HowToChange with existing video datasets on temporal OSC understanding. HowToChange offers an unprecedented scale—with 9.3x more OSC categories and 8.1x more annotated video instances compared to the previous largest collection [49]. Notably, in prior datasets, each state transition is typically coupled with 1 or 2 objects. Consequently, models developed from these datasets are likely limited to recognizing OSCs for the few objects they have been exposed to and unable to generalize to new objects, as we will see in results. In contrast, HowToChange pioneers the open-world formulation, presenting a much broader range of objects associated with state transitions, facilitating the development of models with generalized OSC understanding. Please see Supp. for full data collection and annotation details.

### v2

@@ Abstract
Object State Changes (OSCs) are pivotal for video understanding. While humans can effortlessly generalize OSC understanding from familiar to unknown objects, current approaches are confined to a closed vocabulary. Addressing this gap, we introduce a novel open-world formulation for the video OSC problem. The goal is to temporally localize the three stages of an OSC—the object’s initial state, its transitioning state, and its end state—whether or not the object has been observed during training. Towards this end, we develop V ID OSC, a holistic learning approach that: (1) leverages text and vision-language models for supervisory signals to obviate manually labeling OSC training data, and (2) abstracts fine-grained shared state representations from objects to enhance generalization. Furthermore, we present HowToChange, the first open-world benchmark for video OSC localization, which offers an order of magnitude increase in the label space and annotation volume compared to the best existing benchmark. Experimental results demonstrate the efficacy of our approach, in both traditional closed-world and open-world scenarios. [^1] Project webpage: https://vision.cs.utexas.edu/projects/VidOSC/.
@@ 1 Introduction
However, the development of video OSC is still at a primitive stage. Existing approaches [2, 50, 51, 15, 1, 46, 58] assume a closed vocabulary, where each state transformation is associated with a limited set of objects—often just 1 or 2. For example, the concept of “melting" might be limited to familiar items like butter and marshmallow. Consequently, the learned models are only capable of identifying state changes for objects observed during training and stumble when presented with novel objects. In contrast, in real-world situations a single state transition like “melting" can be linked with a plethora of objects—some ubiquitous and others rare. Importantly, there is an intrinsic connection that threads these OSC processes together. As humans, even if we have never encountered an unusual substance (e.g., jaggery) before, we can still understand that it is experiencing a “melting” process based on its visual transformation cues (see Fig. 1, bottom).
@@
Armed with this new problem formulation, we propose a holistic video OSC learning approach, anchored by two innovative ideas. First, we explore text and vision-language models (VLMs) for supervisory signals during training. We pioneer the use of textual state descriptions to generate a long tail of object state pseudo-labels, by leveraging the remarkable capabilities of VLMs [44, 23, 63] and large language models (LLMs) [53, 54, 40]. This strategy eliminates the need for exhaustive label collection for training data, facilitating large-scale model training. [^2] Note that at test time, our model requires only the video, with no need for additional text, ensuring utmost flexibility and applicability. Second, to confront the open-world challenge, we propose object-agnostic state prediction, consisting of three key techniques: a shared state vocabulary to unify the label space across objects sharing the same state transition (e.g., melting butter and melting marshmallow), temporal modeling to grasp the progression of state changes over time, and object-centric features that better represent the objects during an OSC. These designs equip our model with the ability to generalize state understanding from known objects to novel ones. We term our approach VidOSC.
@@ 2 Related Work
Object State Changes Image-based methods explore the compositional nature of objects and their attributes, including zero-shot recognition of unseen combinations [37, 39, 22, 43, 33, 38, 41, 29], but do not consider the temporal progression of OSCs in video, which brings new challenges. Video-based methods develop human-centric models that leverage state change cues to facilitate action recognition  [15, 58, 1, 46], or explore joint discovery of object states and actions  [2, 50, 51]. Notably, all the existing methods assume a closed world, recognizing only the OSC categories (“known" objects) seen during training. Our approach distinguishes itself in three crucial ways: (1) we introduce a novel open-world333also called “unseen compositions” in object-attribute learning [37, 34]. formulation, where the objects encountered during evaluation can be entirely unseen during training; (2) we adopt an object-centric perspective, allowing for scenarios where OSCs occur with no observable human actions in the video; and (3) we propose to utilize the text modality and VLMs as supervisory signals, greatly scaling up model training and boosting performance.
@@
Vision and Language LLMs [53, 54, 40] have revolutionized various research fields with their exceptional performance across diverse applications. Building on this momentum, the use of web-scale image-text data has emerged as a powerful paradigm in computer vision, with powerful VLMs now advancing an array of image tasks, including zero-shot classification [44], detection [18], segmentation [32] and visual question answering [27, 28]. Similarly, joint video-text representation learning [36, 30, 67, 3] has been advanced by multimodal datasets like HowTo100M [36] and Ego4D [17], which offer large-scale collections of videos paired with their corresponding narrations. These datasets have also facilitated tasks like step discovery [11], localization in procedural activities [35] and long egocentric videos [45]. The vision-language multimodal cycle consistency loss proposed in [14] helps discover long-term temporal dynamics in video. In line with these developments, we propose to leverage the text accompanying instructional videos as well as existing high-performing VLMs to provide supervisory signals that guide the training of our video OSC model and allow learning for the open world.
@@ 3 Approach
Figure 2: Our proposed VidOSC framework: (a) Mining for OSC examples (Sec. 3.3): We leverage ASR transcriptions paired with videos and the capabilities of LLM to automatically mine OSC examples. (b) Pseudo Label Generation (Sec. 3.3): We utilize textual state descriptions and a VLM for supervisory signals during training; (c) Model Training (Sec. 3.2): We develop a video model for object-agnostic state prediction. (d) Model Testing (Sec. 3.1): We propose an open-world formulation, evaluating on both known and novel OSCs. Notably, while we employ the text modality to guide model training, our model is purely video-based and requires no text input at the test phase, ensuring maximum flexibility and applicability. Ground truth for the test set is manually annotated.
@@ 3.1 Video OSC in the Open World
We begin by formally defining an object state change (OSC) in videos as a visually detectable transformation, where an object experiences a change that is not easily reversible, in line with [17]. We characterize an OSC category as an object combined with a specific state transition, such as “chicken + shredding”, and delineate an OSC process through three distinct states: initial (precondition state), transitioning, and end (postcondition state). [^4] Due to real-world data variation, some videos may lack one of the OSC states, and there may be multiple segments in a video corresponding to the same state. Our formulation accounts for all these scenarios. It is essential to highlight that we take an object-centric perspective: the “transitioning state” accounts for instances where the video depicts an active action applied to the object, as well as scenarios where the object undergoes a passive transformation absent of human intervention (e.g., melting, drying).
@@
Next, we propose an open-world problem formulation, capturing the intrinsic long-tail distribution observed in real-world scenarios. Consider $N$ state transitions, each paired with a set of associated objects, denoted by $\mathbf{O}^{n}=\{o_{1}^{n},o_{2}^{n},\cdots,o_{m(n)}^{n}\}$, for $n=\{1,2,\cdots,N\}$, where $m(n)$ denotes the number of objects linked with the $n$-th state transition. Within a specific state transition (like frying), certain objects in the set (such as chicken or onions) are frequently observed, while the combination of the same state transition with other objects (such as cauliflower) appear less often. Motivated by this inherent long-tail, we propose to split $\mathbf{O}^{n}$ into two distinct subsets: $\mathbf{O}_{\text{known}}^{n}$ covering the common combinations observed during training, and $\mathbf{O}_{\text{novel}}^{n}$ comprising the infrequent combinations, which are unseen during training due to their rarity. During inference, the model is evaluated on the entire object set $\mathbf{O}^{n}$ for a comprehensive reflection of the open-world setting.
@@ 3.2 Object-Agnostic State Prediction
Shared State Vocabulary The inherent link among different OSC categories is overlooked in prior work. Common practice involves developing separate models for each OSC category [2, 50] (e.g., one model exclusively for melting butter and another for melting chocolate), or using one single model that still treats every OSC as a distinct entity in the label space [51] (e.g., considering melted butter and melted chocolate as two separate end states). Such a formulation results in an extensive label set with $K=3\times\sum_{n=1}^{N}m(n)$, where 3 denotes the number of OSC states (i.e., initial, transitioning and end) and $\sum_{n=1}^{N}m(n)$ is the total number of OSC categories. This compartmentalized view can inadvertently hinder the model’s generalization ability. Yet, on closer inspection, states across varied objects are intrinsically related. For instance, the “melting” principle remains consistent, even when applied to visually distinct objects like butter and chocolate. This motivates us to group OSC labels that share the same state transition as one, and train a single model per state transition. [^6] See Supp. for the multi-task model variant, where we develop one unified model for all state transitions. By adopting object-agnostic OSC labels, we encourage the model to learn state representations shared by visually different objects, and thus facilitate the transfer from known to novel objects.
@@
Temporal Modeling Recognizing that state changes unfold over time, it is important to capture the temporal dynamics in videos. An object’s state at any frame, be it initial, transitioning, or end, is often best understood in the context of preceding and succeeding frames. Contrary to prior works [2, 50, 51] that rely on isolated frame-wise modeling without considering the temporal progression of video OSCs, we address this gap by proposing a temporally-aware model design. As illustrated in Fig. 2(c), we adopt an encoder-decoder architecture. For the encoding phase, input video features $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{T}\}$ are first projected using an MLP $f_{\text{project}}$, and augmented with sinusoidal positional embeddings, $\mathbf{Z}_{\text{pos}}$. Subsequently, a transformer encoder [56] $f_{\text{transformer}}$ is adopted to capture the temporal dynamics among these features, yielding $\mathbf{Z}=f_{\text{transformer}}(f_{\text{project}}(\mathbf{X})+\mathbf{Z}_{\text{pos}})$. For the decoding phase, a MLP decoder $g$, maps these temporally-aware hidden representations to OSC state predictions: $\tilde{\mathbf{Y}}=g(\mathbf{Z})$. See Sec. 5 for architecture and training details. This design ensures that when predicting the state for a frame, the model has assimilated temporal context from the entire sequence. Essentially, our model exploits the fact that the dynamics of the state change has greater invariance to the object category than how the object looks, emphasizing temporal object transformations over objects’ static appearances, and thereby enhancing generalization to novel OSCs.
@@ 3.3 Text and VLMs as Supervision
Mining for OSC examples Utilizing a vast collection of “how-to" instructional videos as the training source, we develop an automated mining process to capture the rich, real-world variability of OSCs. The motivation is that instructional videos usually have accompanying Automatic Speech Recognition (ASR) transcriptions that offer valuable OSC cues. For example, a speaker mentioning, “so just get your rotisserie chicken and shred it up” suggests that the chicken may undergo a state transition of shredding in the video. Leveraging this fact, we employ LLAMA2 [54] to analyze ASR transcriptions for identifying candidate videos and their associated OSC categories. This text mining stage allows us to corral a long tail of OSCs, discovering likely state change terms—even if rare—from the free-form natural language narrations given by how-to instructors in the training videos. See Fig. 2 (a).
@@
$\hat{y}_{t}=\begin{cases}\text{Background}&\text{if }{\sum}(\mathbf{S}[t,:])<\tau\\ \text{Initial}&\text{elif }\mathbf{S}[t,0]-\mathbf{S}[t,1]>\delta\text{ and }\mathbf{S}[t,0]-\mathbf{S}[t,2]>\delta\\ \text{Transitioning}&\text{elif }\mathbf{S}[t,1]-\mathbf{S}[t,0]>\delta\text{ and }\mathbf{S}[t,1]-\mathbf{S}[t,2]>\delta\\ \text{End}&\text{elif }\mathbf{S}[t,2]-\mathbf{S}[t,0]>\delta\text{ and }\mathbf{S}[t,2]-\mathbf{S}[t,1]>\delta\\ \text{Ambiguous}&\text{otherwise}\end{cases}$
@@
To further refine these labels, we enforce a causal ordering constraint. Given the inherent progression of OSCs, the anticipated order is initial states followed by transitioning states, and finally, the end states. Any frame whose pseudo label does not respect this natural progression is re-assigned to the ambiguous category. Our training employs a cross-entropy loss between pseudo label $\hat{y}_{t}$ and the corresponding model prediction $\tilde{y}_{t}$, with ambiguous frames excluded to maintain clarity and distinction among the state labels. See Section 4.2 in Supp. for detailed pseudo label analysis.
@@ 4 The HowToChange Dataset
Existing OSC datasets fall short in capturing the open-world’s diversity and long-tail distribution of OSCs. To address this gap, we introduce HowToChange. It encompasses varied state transitions coupled with a diverse range of objects, providing an authentic reflection of their real-world frequency—from the commonplace to the rare.
@@
Ground Truth Label Collection (Evaluation Set) To facilitate thorough evaluation, we obtain manual annotations for a subset of 5,424 videos from the collected dataset. [^7] Our training is purely guided by a VLM and requires no ground truth labels. The annotations are reserved exclusively for evaluation. The annotation workflow is as follows: each annotator is presented with a video segment along with an OSC category that was previously identified by LLAMA2. The annotator has the option to reject the video segment if it does not contain the specified OSC. Otherwise, they label the time ranges corresponding to the initial, transitioning, and end states of the OSC. Adhering to our object-centric emphasis, annotators are instructed to label based on the visual changes of the object, rather than human-centric actions, and exclude time ranges where the object of interest is not visible, ensuring clean and focused temporal labels. Fig. 3 provides the distribution of annotated videos. On average, we collect 271 annotations per state transition, with a video duration of 41.5 seconds, and 12.9% of videos belong to the novel category. The entire annotation process required around 1,507 hours by 30 professional annotators.
 Datasets # Obj # ST # OSC ObjPer # Videos GT Label?
@@
Table 1: Comparison with existing video datasets focusing on object states. ‘Obj’ and ‘ST’ represent objects and state transitions, respectively. ‘ObjPer’ denotes the average number of objects associated with each state transition; higher values indicate more need to generalization across objects. We present the first open-world benchmark for temporal video OSC, with an order of magnitude increase in OSC vocabulary and annotation volume.
Dataset Comparison Table 1 compares HowToChange with existing video datasets on temporal OSC understanding. HowToChange offers an unprecedented scale—with 9.3x more OSC categories and 8.1x more annotated video instances compared to the previous largest collection [50]. Furthermore, notably, in prior datasets, each state transition is typically coupled with 1 or 2 objects, preventing subsequent models from generalizing to new objects, as we will see in results. In contrast, HowToChange pioneers the open-world formulation, presenting a much broader range of objects associated with each state transition—from 6 to 55 objects per state transition, and averaging 20. This facilitates the development of models with generalized OSC understanding. Please see Supp. for full data collection and annotation details.